# FoMo: A Foundation Model for Mobile Traffic Forecasting with Diffusion Model

### v1

@@ 3.3.2. Self-supervised masked diffusion model
This part reconstructs the masked mobile token $E_{x}(X)$ that includes the following three important processes.
$\bullet$ Forward noising process. After completing the multi-task masking process, the original mobile token $E_{x}(X)$ is divided into two parts: the masked portion $e$ requiring to be reconstructed, and the unmasked observation $o$:
@@
Afterward, $e_{k}$ is fed into the transformer-based denoising network for further feature extraction.
$\bullet$ Adaptive conditioning mechanism. To fully capture the dependencies between conditions and network traffic latent features, we employ an adaptive layernorm method (Peebles and Xie, 2023). The method reshapes the scale and shift parameters of the layernorm of transformers by referring to the given conditions, which is proven to offer better effectiveness and computational efficiency (Perez et al., 2018). It can be formulated as follows:
@@
where $\mathcal{F}_{\theta}$ and $\mathcal{A}_{\theta}$ denote linear layer and attention layer. $\alpha,\beta,\gamma$ are residual, scale, and shift parameters, respectively.
$\bullet$ Masked loss function. The denoising network aims to fit the posterior distribution of the diffusion process to predict the mean noise, ultimately reconstructing the final network traffic through an output decoder. Our objective thus emphasizes the reconstruction accuracy of the masked portion, which can be given as indicated in (9):
@@ 3.4. Urban context-aware fine-tuning
Mobile traffic is not only a spatio-temporal sequence but also influenced by urban contexts. Extensive literature has demonstrated that incorporating contextual features can enhance traffic prediction performance (Wang et al., 2023; Yu et al., 2021; Xu et al., 2022). Therefore, we propose a fine-tuning scheme, where two modules are encapsulated into the pre-trained model:
$\bullet$ Contextual integration: The module introduces mobile users and POI distributions, which aims to integrate human dynamics with the static features of urban environments.
[…]
3.4.1. Contextual integration
Mobile user refers to the number of users accessing the network. Similar to mobile traffic, it is inherently a spatio-temporal sequence that is denoted as $U\in\mathbb{R}^{H\times V\times T}$. We apply the same processing method as traffic tokens where we perform tokenization on mobile users as $c^{u}\in\mathbb{R}^{(H^{\prime}\times V^{\prime}\times T^{\prime})\times(h_{0}\times v_{0}\times t_{0})}$, allowing this data to be directly input into the network for training.
Urban POIs reflect the static distribution of urban layout that can be denoted as $P\in\mathbb{R}^{H\times V}$. We cannot directly utilize spatio-temporal tokenization methods. Hence, we design a dynamic POI transformation scheme. Although the distribution of POIs is static, the impact of different categories of POIs on human behavior varies across different times, leading to corresponding variations in mobile traffic. For example, restaurant-type POIs typically show higher traffic during lunchtime and evening. We first extract the intrinsic static features of POI distribution, which can be written as:
@@ 3.4.1. Contextual integration
where $\sigma$ is the Sigmoid activation function, $W^{s}$ and $B^{s}$ are the weight and bias parameters of a linear layer. Inspired by (Wang et al., 2017a), we introduce temporal projection $\tau(t)$ related to timestamp (we use an MLP network). Then, the indicators and static features of POIs $h^{s}_{p}$ are concatenated as $h^{(0)}=h^{s}_{p}\oplus\tau(t)$, where $\oplus$ denotes vector concatenation. We utilize an MLP network to fuse both the time embeddings and POI features:
(13) $h^{(l+1)}=\sigma(W^{l}\cdot h^{(l)}+B^{l}),\ \ \ l\in[0,L),$
where $h^{(l)}$ is the hidden feature of the $l$-th layer of MLP, $W^{l}$ and $B^{l}$ is the weight and bias parameters in each layer. In this way, we can obtain spatio-temporal dynamic representations as $h_{p}^{d}=h^{(L)}\in\mathbb{R}^{H\times V\times T}$. The final features of POI can be calculated via the same tokenization method as $c^{p}\in\mathbb{R}^{(H^{\prime}\times V^{\prime}\times T^{\prime})\times(h_{0}\times v_{0}\times t_{0})}$. The ultimate contextual feature tokens can be denoted as $y=c^{u}+c^{p}$.
@@ 3.4.2. Context-aware alignment
Contextual features can be viewed as another modality of data that contains much information related to mobile traffic. We propose a contrastive learning algorithm (He et al., 2020; Chen et al., 2020a) to establish bridges between mobile traffic and contextual features.
$\bullet$ Contrastive sample construction. We define positive samples as mobile traffic tokens and contextual feature tokens within the same spatio-temporal block, denoted as $(e,y)$; while negative samples are defined as the two types of tokens from different spatio-temporal blocks, denoted $(e^{\prime},y)$.
$\bullet$ Contrastive goal. Our goal is to maximize the mutual information between the traffic feature $e$ and contextual feature $y$. According to previous research (van den Oord et al., 2019), a density ratio can be utilized while preserving the mutual information as $I(e,y)\propto\frac{p(e|c)}{p(e)}$, and the maximization problem is equivalent to minimizing the InfoNCE loss that yields:
(14) $L=\underset{e\in\mathbb{B}}{\mathbb{E}}-log\frac{I(e,y)}{I(e,y)+\sum_{e^{\prime}}I(e^{\prime},y)}\leq log(N)-I(e,y),$
@@
$\bullet$ Contrastive learning with diffusion paradigm. For diffusion models, the objective is to fit the posterior distribution through a neural network $p_{\theta}(x_{k-1}|x_{k})$ that maximizes the likelihood of $p_{\theta}(x_{0})$. We clarify that the objective is fundamentally equivalent to that of InfoNCE in contrastive learning. We use $p_{\theta}$ to represent the probability in the mutual information as $I(e,y)=p_{\theta}(e_{0:K}|y)/p_{\theta}(e_{0:K})$. In this way, the origin InfoNCE loss can be rewritten as:
where $e^{\prime}$ denotes all negative samples. Referencing the parameterization $p_{\theta}(x_{k-1}|x_{k})=\sum_{x_{0}\in q}q(x_{k-1|x_{k},x_{0}})p_{\theta}(x_{0}|x_{k})$ in  (Zhu et al., 2023; Austin et al., 2021), the above loss can be further formulated as:
$\ \ =L^{e}_{vb}-logN\sum\nolimits_{e^{\prime}}L^{e^{\prime}}_{vb}$
where symbol $\doteq$ denotes the loss function we used during the model training process and $\lambda$ is a scaling parameter proportional to $logN$. Building on the above analysis, we can guide FoMo’s fine-tuning strategy:
By optimizing the Mean Squared Error (MSE) of positive and negative samples, we can achieve alignment between mobile traffic and contextual information features.
$\bullet$ Partially frozen fine-tuning. During the fine-tuning process, we froze the main parameters of the pre-trained model, including the attention layer, linear layer, and MLP network, to preserve the model’s ability to learn general spatio-temporal features of mobile traffic. We primarily update the parameters in the adaptive layernorm, contextual integration module, and output decoder layer. By partially updating these components, the time and computational cost of the fine-tuning process can be reduced (Shen et al., 2021; Ye et al., 2023).
@@ 4. Evaluation
We perform evaluations on nine real-world datasets to discuss the universality of the FoMo with 13 baselines. The experiments need to address the following three questions.
  • RQ1: How does it perform in multi-task forecasting across multiple datasets?
  • RQ2: How does the model perform in zero-shot and few-shot learning tasks?
  • RQ3: How does the proposed model perform in terms of scalability?
4.1. Evaluation settings
@@ 4.1.1. Datasets
Mobile traffic data. We collect mobile traffic data from 7 cities of varying scales in China, which encompass downlink traffic including both 4G and 5G data. The time granularity of the data ranges from 15 minutes to 1 hour. Additionally, we utilized mobile traffic data from 2 other cities in China and Germany to validate FoMo’s zero/few-shot capabilities, where the Germany dataset is generated using the traffic prediction method described in (Xu et al., 2021).
Table 1. Description of datasets
Dataset Usage Data Mobile Mobile Time
description traffic users granularity
Beijing 5G data, October, 2021, 4000+ BSs $\checkmark$ $\checkmark$ 30 min
Shanghai 4G data, August, 2014, 5000+ BSs $\checkmark$ $\checkmark$ 30 min
Nanjing Model training 5G data, February to March, 2021, 6000+ BSs $\checkmark$ $\checkmark$ 15 min
Nanchang 5G data, July, 2023, 5000+ BSs $\checkmark$ $\checkmark$ 30 min
Munica shot tests 4G generated data, 2022 $\checkmark$ 1 hour
Urban Contextual data. We collect mobile user data in conjunction with mobile traffic data in each dataset. We crawl POI data from each city through public map services, including 15 categories related to living, entertainment, and other aspects. We summarize the collected data in Table 1.
Table 2. Performance of short-term prediction task. Bold numbers denote the best results and $\underline{underline}$ numbers denote the second-best results. The proposed FoMo has the best prediction performance across 7 datasets.
@@ 4.1.2. Baselines
We select a total of 13 baselines across 4 major types.
$\bullet$ Statistical models. Historical moving average method (HA) and ARIMA method that integrate autoregressive with acreage moxing (Xu et al., 2023).
$\bullet$ Natural language-based model. Time-LLM (Jin et al., 2024) describes time series features using natural language and uses these descriptions as prompts into a natural language pre-trained model (LLAMA-7B) for forecasting. Tempo (Cao et al., 2024) designs temporal prompts with trend and seasonal features for pre-trained models (GPT-2) to predict time series.
$\bullet$ Spatio-temporal-based models. They primarily forecast mobile traffic as spatio-temporal series via autoregression, decomposition, and spatial convolution. TimeGPT (Garza and Mergenthaler-Canseco, 2023) replaces the Feedforward layer in the transformer with a CNN network and is trained on vast spatio-temporal data. Laglla- ma (Rasul et al., 2024) uses a set of lag indices to capture different periodic correlations in the time series. PatchTST  (Nie et al., 2023) decomposes time series into multiple segments and uses transformers for feature extraction. UniST (Yuan et al., 2024a) segments spatio-temporal data and fine-tunes the model using features like geographical proximity and temporal correlations.
$\bullet$ Dedicated models for mobile traffic forecasting. In addition to the general spatio-temporal forecasting models, we select 5 models dedicated to mobile networks. SpectraGAN (Xu et al., 2021) converts mobile traffic generation into an image generation problem and utilizes a CNN-based GAN network for traffic forecasting. KEGAN (Hui et al., 2023) is a hierarchical GAN that utilizes a self-constructed Urban Knowledge Graph (UKG) to explicitly incorporate urban features during the forecasting process. ADAPTIVE (Zhang et al., 2023) leverages the UKG and a BS aligning scheme to transfer mobile traffic knowledge from one city to another. CSDI (Tashiro et al., 2021) is a conditional diffusion model that uses a masking method for mobile traffic forecasting. KG-Diff (Chai et al., 2024b) inputs both the seasonal part of mobile traffic and the UKG into a diffusion model, aiding the model in understanding the spatio-temporal dynamics.
4.1.3. Metrics
We choose 3 metrics to investigate the performance of the algorithms. MAE evaluates the similarity of generated values $S$ and real values $\hat{S}$, which can be expressed as $e=Avg(|S-\hat{S}|)$. RMSE measures the average magnitude of the errors between predicted values and actual observed values, which can be expressed as $e=\sqrt{Avg(|S-\hat{S}|^{2})}$. Jensen–Shannon divergence (JSD) is a commonly used metric to measure the similarity between two distributions.
@@ 4.2. Multitask forecasting (RQ1)
Table 3. Performance of Long-term prediction task. Bold numbers denote the best results and $\underline{underline}$ numbers denote the second-best results. The proposed FoMo has the best prediction performance across 7 datasets.
@@
Lagllama 0.2318 0.1879 0.1453 0.08742 $\underline{0.0960}$ $\underline{0.0683}$ $\underline{0.1115}$ $\underline{0.0959}$ $\underline{0.1091}$ $\underline{0.0788}$ 0.1684 0.0889 0.1076 0.0439
@@
FoMo (our) 0.1035 0.0696 0.0983 0.0679 0.0818 0.0532 0.0849 0.0570 0.0853 0.0576 0.1206 0.0563 0.0518 0.0197
In our experiments, the temporal length is 64. For short-term prediction, the model forecasts 16 future points using the previous 48. For long-term prediction, the model forecasts 48 future points using the previous 16. For data generation, the model predicts all 64 points based on the current timestamp.
4.2.1. Short-term prediction
The results of the short-term forecasting task are presented in Table 2. Since sufficient historical data is available for reference, most baselines, leveraging their temporal feature extraction modules, effectively predict short-term changes. Our proposed FoMo outperforms the best baseline across multiple datasets, where FoMo can improve the RMSE metric by up to 29.1% (Shandong dataset) and the MAE metric by up to 50% (Nanjing-4G dataset), which exhibits stronger generalization capabilities compared to other models. Through the adaptive layernorm module, the diffusion model iteratively integrates contextual features and leverages the transformer to capture long-term dependencies between mobile traffic and the environment. We believe this correlation can transfer across different cities, improving the model’s generalization capability.
4.2.2. Long-term prediction
The long-term forecasting results are also shown in Table 3. For this task, the lack of sufficient historical observations often leads to performance degradation in some baselines. However, FoMo consistently achieves the best performance, it achieves a maximum improvement of 31.4% in RMSE (Beijing dataset) and a maximum enhancement of 40.6% in MAE (Nanjing-4G dataset), which showcasing its adaptability to various tasks.
4.2.3. Mobile traffic generation
Table 4. Performance of generation task. Bold numbers denote the best results and $\underline{underline}$ numbers denote the second-best results. The proposed FoMo has the best generation performance across 7 datasets.
@@ 4.2.3. Mobile traffic generation
CSDI 0.3385 0.1431 $\underline{0.2331}$ $\underline{0.1025}$ 0.4044 0.1844 0.3875 0.2516 0.3416 0.2163 0.2895 0.1734 0.2666 0.2170
KG-Diff $\underline{0.2397}$ $\underline{0.1299}$ 0.2504 01245 $\underline{0.2223}$ $\underline{0.1214}$ $\underline{0.2323}$ $\underline{0.1514}$ 0.2683 0.1356 $\underline{0.2744}$ $\underline{0.1363}$ $\underline{0.2041}$ $\underline{0.1101}$
FoMo (our) 0.2013 0.0894 0.2259 0.1002 0.1971 0.0948 0.2164 0.0938 0.2226 0.1043 0.2494 0.1159 0.1558 0.0993
As shown in Table 4, the absence of historical observation of the generation task prevents some existing baselines from completing the task. Nevertheless, the FoMo still achieved strong generative results. It can achieve up to a 23.7% improvement in the JSD metric (Shandong dataset), and a maximum enhancement of 38.1% in MAE (Nanjing-4G dataset). This is due to the contextual feature fusion module used during fine-tuning, which captures the correlation between contextual and mobile traffic features through contrastive learning. This allows the model to infer potential traffic distribution based on environmental changes, even without historical data.
4.2.4. Visualization
To intuitively demonstrate our model’s universality for different tasks, we select two datasets as examples (Beijing and Nanchang) and plot the forecasting results in Figure LABEL:visual_res. From left to right, it represents the tasks of short-term prediction $\rightarrow$ long-term prediction $\rightarrow$ traffic generation. The blue-shaded area indicates the model’s predicted results, while the unshaded area represents historical observations. FoMo generates mobile traffic closely aligned with real values across all tasks, accurately predicting periodic trends and capturing fast dynamics, which shows that our FoMo model achieves forecasting across multiple cities and tasks, highlighting its generalization capability.

### v3

@@ 3.2.3. Self-supervised masked diffusion model
After completing the multi-task masking process, the original mobile token $E_{x}(X)$ is divided into two parts: the masked portion $e$ requiring to be reconstructed, and the unmasked observation $o$:
@@
Afterward, $e_{k}$ is fed into the transformer-based denoising network for further feature extraction. To fully capture the dependencies between conditional observation and mobile traffic, we employ an adaptive conditioning method (10377858, 40). The method reshapes the scale and shift parameters of the layernorm of transformers by referring to the given conditions, which is proven to offer better effectiveness and computational efficiency (10.5555/3504035.3504518, 41). It can be formulated as follows:
@@
where $\mathcal{F}_{\theta}$ and $\mathcal{A}_{\theta}$ denote linear layer and attention layer. $\alpha,\beta,\gamma$ are residual, scale, and shift parameters, respectively. The denoising network aims to fit the posterior distribution of the diffusion process to predict the mean noise, ultimately reconstructing the final network traffic through an output decoder. Our objective thus emphasizes the reconstruction accuracy of the masked portion, which can be given as indicated in (5):
@@ 3.3. Urban context-aware fine-tuning
Mobile traffic is not only a spatio-temporal sequence but also influenced by urban contexts. We thereby propose an urban context-aware fine-tuning scheme that integrates human dynamics and POIs into the UoMo, as shown in Fig 3.
3.3.1. Contextual data transformation
Mobile user refers to the number of users accessing the network, which can fully characterize the human dynamics in mobile networks. Similar to mobile traffic, it is inherently a spatio-temporal sequence and denoted as $U\in\mathbb{R}^{H\times V\times T}$. We apply the same processing method as traffic tokens where we perform tokenization on mobile users as $c^{u}\in\mathbb{R}^{(H^{\prime}\times V^{\prime}\times T^{\prime})\times(h_{0}\times v_{0}\times t_{0})}$, allowing this data to be directly input into the network for training.
POIs reflect the static distribution of urban layout. For each base station’s coverage area, we count the total number of each POI category within the coverage area so that we can obtain a POI vector $P\in\mathbb{R}^{H\times V}$. Although the distribution of POIs is static, the impact of different categories of POIs on human behavior varies across different times, leading to corresponding variations in mobile traffic. For example, restaurant-type POIs typically show higher traffic during lunchtime and evening. In this regard, we design a dynamic POI transformation scheme. We first extract the intrinsic static features of POI distribution, which can be written as:
@@ 3.3.1. Contextual data transformation
where $\sigma$ is the Sigmoid activation function, $W^{s}$ and $B^{s}$ are the weight and bias parameters of MLP network. We further utilize an MLP network $\tau(t)$ to project timestamp as temporal embeddings, where we use the nn.Embedding layer to encode the 2D-vector $t=[day,hour]$ and fuse the two embeddings using an MLP network, then we fuse the static POI feature with the temporal indicators:
(9) $h^{d}_{p}=\sigma(W^{l}\cdot[h^{s}_{p}\oplus\tau(t)]+B^{l}),$
where $\oplus$ denotes vector concatenation, $W^{l}$ and $B^{l}$ are the learnable parameters. In this way, we can obtain spatio-temporal dynamic representations as $h_{p}^{d}\in\mathbb{R}^{H\times V\times T}$. The final features of POI can be calculated via the same tokenization method as mobile traffic data: $c^{p}\in\mathbb{R}^{(H^{\prime}\times V^{\prime}\times T^{\prime})\times(h_{0}\times v_{0}\times t_{0})}$. The ultimate contextual feature tokens can be denoted as $y=c^{u}+c^{p}$.
@@ 3.3.2. Context-aware alignment
To establish bridges between mobile traffic and contextual features, we propose a contrastive learning algorithm. We define positive samples as mobile traffic tokens and contextual feature tokens within the same spatio-temporal block, denoted as $(e,y)$; while negative samples are defined as the two types of tokens from different spatio-temporal blocks, denoted $(e^{\prime},y)$. Our goal is to maximize the mutual information between the traffic feature $e$ and contextual feature $y$. According to previous research (oord2019representationlearningcontrastivepredictive, 49), a density ratio can be utilized while preserving the mutual information as $I(e,y)\propto\frac{p(e|c)}{p(e)}$, and the maximization problem is equivalent to minimizing the InfoNCE loss that yields:
(10) $min-\underset{e\in\mathbb{B}}{\mathbb{E}}log\frac{\frac{p(e|c)}{p(e)}}{\frac{p(e|c)}{p(e)}+\sum_{e^{\prime}}\frac{p(e^{\prime}|c)}{p(e^{\prime})}}\geq log(N)-I(e,y),$
@@
We claim in Lemma 1 that training the diffusion model with positive and negative samples is equivalent to minimizing the InfoNCE loss in contrastive learning of Eq (10):
Lemma 1: By optimizing the Mean Squared Error (MSE) of positive and negative samples via Eq (11), we can achieve alignment between mobile traffic and contextual features.
(11) $L\approx\mathbb{E}\bigg{\{}\bigg{(}\|\epsilon-\epsilon_{\theta}(e,k|y)\|^{2}-\lambda\sum\nolimits_{e^{\prime}}\|\epsilon-\epsilon_{\theta}(e^{\prime},k|y)\|^{2}\bigg{)}\odot m\bigg{\}}.$
The proof of the lemma is provided in the appendix A.3. During the fine-tuning process, we partially froze the main parameters of the pre-trained model, including the attention layer, linear layer, and MLP network, to preserve the model’s ability to learn general spatio-temporal features of mobile traffic. We primarily update the parameters in the adaptive conditioning and output decoder layers. By partially updating these components, the time and computational cost of the fine-tuning process can be reduced.
@@ 4. Evaluation
We perform evaluations on 9 real-world datasets to evaluate the UoMo with 13 baselines. The evaluations need to address the following 2 questions.
$\bullet$ RQ1: How does it perform in multi-task forecasting?
$\bullet$ RQ2: How does it perform in zero-shot and few-shot learning?
4.1. Evaluation setting from live mobile system
@@ 4.1.1. Datasets
Mobile traffic data. We collect mobile traffic data of live networks from 7 cities of varying scales in China, encompassing downlink traffic including 4G and 5G data. The time granularity of the data ranges from 15 minutes to 1 hour. Additionally, we utilize mobile traffic data from 2 other cities in China and Germany to validate UoMo’s zero/few-shot capabilities.
Urban Contextual data. We collect mobile user data and mobile traffic data in each dataset. We crawl POI data from each city through public map services, including 15 categories related to living, entertainment, and other aspects.
Table 1. Performance of short-term prediction task. Bold numbers denote the best results and $\underline{underline}$ numbers denote the second-best results. The proposed UoMo has the best prediction performance across 7 datasets.
@@ 4.1.2. Baselines
We select 13 baselines with 4 types. i). Statistical models. Historical moving average (HA) and ARIMA (10.1145/3573900.3591123, 59). ii) NLP-based model. Time-LLM (jin2024timellmtimeseriesforecasting, 23) and Tempo (cao2024tempopromtrained, 7) describes time series features using natural language and uses these descriptions for forecasting. iii). Spatio-temporal based models. CSDI (NEURIPS2021_cfe8504b, 47), TimeGPT (garza2023timegpt1, 15), Lagllama (rasul2024lagllamafoundationmodelsprobabilistic, 43), PatchTST (nie2023timeseriesworth64, 38), and UniST (10.1145/3637528.3671662, 69) forecast mobile traffic as spatio-temporal series via autoregression, decomposition, and spatial convolution. iv). Dedicated models for mobile networks. SpectraGAN (10.1145/3485983.3494844, 58), KEGAN (10.1145/3580305.3599853, 21), ADAPTIVE (10.1145/3580305.3599801, 71), and Open-Diff (10.1145/3637528.3671544, 8) utilizes dedicated contextual data to generate mobile traffic. We describe the collected data and baselines in the appendix A.1 and A.2.
@@ 4.2. Multitask forecasting (RQ1)
In our experiments, the temporal length is 64. For short-term prediction, the model forecasts 16 future points using the previous 48. For long-term prediction, the model forecasts 48 future points using the previous 16. For data generation, the model predicts all 64 points based on the current timestamp. We primarily use the 35M UoMo model (with 16 transformer layers and a hidden feature size of 256), and the other scaling evaluations are provided in Table 6.
To intuitively demonstrate our model’s universality for different tasks, we select two datasets as examples (Beijing and Nanchang) and plot the forecasting results in Figure LABEL:visual_res. From left to right, it represents the tasks of short-term prediction $\rightarrow$ long-term prediction $\rightarrow$ traffic generation. The blue-shaded area indicates the model’s predicted results, while the unshaded area represents historical observations. UoMo generates mobile traffic closely aligned with real values across all tasks, accurately predicting periodic trends and capturing fast dynamics, which shows that our UoMo model achieves forecasting across multiple cities and tasks, highlighting its generalization capability.
Table 2. Performance of Long-term prediction task.
@@
Lagllama 0.2318 0.1879 0.1453 0.0874 $\underline{0.0960}$ 0.0683 $\underline{0.1115}$ 0.0959 $\underline{0.1091}$ 0.0788 0.1684 0.0889 0.1076 0.0439
@@
UoMo (our) 0.1035 0.0696 0.0983 0.0679 0.0818 0.0532 0.0849 0.0570 0.0853 0.0576 0.1206 0.0563 0.0518 0.0197
Improvement 27.41% 31.36% 11.44% 11.35% 14.79% 10.14% 23.85% 10.09% 21.81% 18.76% 16.54% 26.21% 16.72% 22.13%
Table 3. Performance of generation task.
@@
CSDI 0.3385 0.1431 0.2331 0.1025 0.4044 0.1844 0.3875 0.2516 0.3416 0.2163 0.2895 0.1734 0.2666 0.2170
Open-Diff $\underline{0.2155}$ $\underline{0.1112}$ $\underline{0.2299}$ $\underline{0.1020}$ $\underline{0.2114}$ $\underline{0.1149}$ $\underline{0.2275}$ $\underline{0.1322}$ 0.2296 0.1326 $\underline{0.2624}$ $\underline{0.1222}$ $\underline{0.1977}$ $\underline{0.1203}$
UoMo (our) 0.2013 0.0894 0.2259 0.1002 0.1971 0.0948 0.2164 0.0938 0.2226 0.1043 0.2494 0.1159 0.1558 0.0993
Improvement 6.58% 19.60% 1.74% 1.76% 6.76% 17.49% 4.87% 29.04% 2.58% 19.02% 4.95% 5.15% 21.19% 17.45%
$\bullet$ Short-term prediction. The results are presented in Table 1. Since sufficient historical data is available for reference, most baselines, leveraging their temporal feature extraction modules, effectively predict short-term changes. Overall, UoMo improved the average RMSE performance by 17.80% and the average MAE by 27.85% across 7 datasets, where UoMo can improve the RMSE metric by up to 29.1% (Nanchang-4G dataset) and the MAE metric by up to 50% (Nanjing-4G dataset), which exhibits stronger generalization capabilities compared to other models. Through the adaptive layernorm module, the diffusion model iteratively integrates contextual features and leverages the transformer to capture long-term dependencies between mobile traffic and the environment. We believe this correlation can transfer across different cities, improving the model’s generalization capability.
$\bullet$ Long-term prediction. The results are also shown in Table 2. For this task, the lack of sufficient historical observations often leads to performance degradation in some baselines. However, UoMo consistently achieves the best performance, It improves the average RMSE performance by 18.93% and the average MAE by 18.57% across 7 real-world datasets, with a maximum improvement of 31.36% in MAE and a maximum enhancement of 27.41% in RMSE (Beijing dataset), which showcases its adaptability to various tasks.
$\bullet$ Mobile traffic generation. As shown in Table 3, the absence of historical observation of the generation task prevents some existing baselines from completing the task. Nevertheless, the UoMo still achieved strong generative results. It achieves a 6.95% improvement in the average JSD performance and a 15.6% improvement in the average MAE across all the datasets, with up to 21.19% improvement in the JSD metric (Shandong dataset), and a maximum enhancement of 29.04% in MAE (Nanjing-4G dataset). This is due to the contextual feature fusion module used during fine-tuning, which captures the correlation between contextual and mobile traffic features through contrastive learning. This allows the model to infer potential traffic distribution based on environmental changes, even without historical data.
Table 4. Zero/Few-shot evaluation on FY and NN datasets.
Model
Zero-shot
5% Few-shot
10% Few-shot
UoMo (FY) 0.156 0.034 0.023
LagLlama (FY) 0.162 0.094 0.053
UoMo (NN) 0.134 0.051 0.021
LagLlama (NN) 0.152 0.088 0.026